OpenStudy (anonymous):

Help ASAP I WILL FAN & MEDAL!! x + 12.85 = 0 (Points : 2) 0 1 12.85 -12.85

OpenStudy (igreen):

x + 12.85 = 0 You have to subtract 12.85 to both sides, so what's 0 - 12.85?

OpenStudy (anonymous):

Isn't it -12.85

OpenStudy (anonymous):

0

OpenStudy (igreen):

Yep, you got it! So that's your answer.

OpenStudy (anonymous):

Simplifying x + 12.85 = 0 Reorder the terms: 12.85 + x = 0 Solving 12.85 + x = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-12.85' to each side of the equation. 12.85 + -12.85 + x = 0 + -12.85 Combine like terms: 12.85 + -12.85 = 0.00 0.00 + x = 0 + -12.85 x = 0 + -12.85 Combine like terms: 0 + -12.85 = -12.85 x = -12.85 Simplifying x = -12.85

OpenStudy (anonymous):

This is what a site told me @iGreen

OpenStudy (anonymous):

So is it 0 or -12.85 ??

OpenStudy (igreen):

It's -12.85..at the end it says x = -12.85

OpenStudy (anonymous):

Okay, thank you :)

OpenStudy (igreen):

No problem.